OpenStudy (anonymous):

in a 20-row theater,the number of seats in a row increases by three with each successive row.the first row has 18 seats. a:write an arithmetic series to represent the number of seats in the theater. please help im stuck!

OpenStudy (anonymous):

This is easy. Make a chart, it should be very clear to you after you see a chart of what you are trying to look at.

OpenStudy (anonymous):

would i not need a formula?

OpenStudy (anonymous):

There is a way, but I find it more clear to see by using a chart.

OpenStudy (anonymous):

i dont know how to even set up the equation

OpenStudy (anonymous):

Make a chart, it will help you much more.

OpenStudy (jdoe0001):

@catrina so, I assume you've covered arithmetic sequences?

OpenStudy (anonymous):

how do i set the chart up?as : 1st row-18 (+3) 2nd-21(+3) ... 20th row-75

OpenStudy (anonymous):

Yes, that is correct.

OpenStudy (anonymous):

im helping my boyfriend i have already done this but it was 2 years ago and i cant remember a thing

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

ok now how do i set up the equation to show the number of seats?

OpenStudy (jdoe0001):

well, the sum of a partial sequence is \(\bf \Large s_n = \cfrac{n}{2}(a_1+a_n)\)

OpenStudy (anonymous):

so thts the formula?

OpenStudy (jdoe0001):

for the sum from the 1st term up to the Nth term, yes

OpenStudy (jdoe0001):

hmm, but you're not being asked on the equation itself, so much as to write the sequence itself, I think

OpenStudy (anonymous):

ok so basicly i wud set it up as S20=20/2(18+20)?

OpenStudy (jdoe0001):

well, the 1st term is 18, the 20th term is not 20 though but I gather is not what's being asked what's being asked is, to "write an arithmetic serie" not to give the total amount of seats in the theater

OpenStudy (anonymous):

yea but it says write one to represent the total amount

OpenStudy (jdoe0001):

well, well what's the common ratio in the sequence you think?

OpenStudy (anonymous):

it says 3

OpenStudy (jdoe0001):

so to find the 20th term, the Nth term, n = 20, \(\bf \Large a_n = a_1+(n-1)d\) n = the nth term d = common ratio

OpenStudy (anonymous):

ok,so a20=18+(20-1)3 ?

OpenStudy (jdoe0001):

hmm, i notice up above you already had it as 75, so , just use the Sum equation then :)

OpenStudy (jdoe0001):

"catrina ok so basicly i wud set it up as S20=20/2(18+20)?" yes

OpenStudy (anonymous):

ok thnx